There have been many reports of trouble with eDocs today, from both on and off campus. Tulio is currently working on the situation.
Some student e-mail is currently down due to a failure of one of the Exchange servers. See this link for more info.
Also, one of the old VMware servers has also suffered multiple drive failures. Some servers hosted on that system are no longer responding.
The Network Services team is working on each of these issues and well keep you updated as the day progresses.
The Camtasia Relay server is UP and working normally.
On Sunday afternoon and then again on Monday morning users were receiving an http 500 error. This error was caused by a disconnect of the large network storage drive D, where all the presentations are located.
The D drive is now connected to the SAN and System Support is investigating the cause of the iscsi initiator disconnect.
Troy Fugnitti joins the Network Services team today. Troy will be providing support for the Computer Science VMware cluster and working with the other Network Services team members supporting servers, applications, and infrastructure. He will normally be working an afternoon/evening shift during the week, so he will be available on site to help resolve issues when many of the rest of us have already headed home. Troy’s phone extension is 68576.
Please welcome Troy next time you see him.
We have 6 Access points in Lincoln Res Hall that did not restart after the power outage on Saturday. Four units on the third floor and two on the west wing of the second floor are not functioning properly. We are working to restore signal.
At 4:27pm on 11/26/11 we received an email from Emily Boles indicating that Blackboard application had a Tomcat error message.
After troubleshooting we found out that the BB database server was down and not listening. After starting the Oracle instance and listener and restarting the Blackboard services on all the web-app servers, the Blackboard application is back up. We checked the uptime of the database server, and it looks like the Blackboard Oracle database server was rebooted (~3:40pm) and the instance did not start correctly after the reboot. My guess is that this problem was due to the power outage.
This was sent to all faculty and online faculty:
I am happy to share the good news that ITS has purchased 3 additional processing licenses for Camtasia Relay – which means we can process more recordings and do it more quickly than ever before.
The Camtasia Relay server will be down on Tuesday, November 22 from 7am-Noon to apply these licenses and to upgrade the server hardware. There is no need for you to do anything; Camtasia Relay will look and function as normal, except for faster processing.
The October 2011 Helpdesk contact report can be found at this website, courtesy of Jeff (thanks jeff!)
Systems Support will be migrating the following servers on the morning of Thursday, November 17, during the normal maintenance window (between 5AM and 7AM):
Once completed, each effected user/group will be contacted.
When- Nov 22, 2011 9:00 am
What- Relocation of the campus Secondary Distribution Switch
Why- To improve the security and environmental controls of the switch, to simplify the campus network topology, and to prepare for virtualization of the two distribution switches.
Impact- Some wireless access points will be unavailable during the move. Some locations will experience up to a 45 second delay while network topology changes.
Duration- We anticipate the move and reactivation of the switch to take 2-3 hours.
The secondary distribution switch and the secondary ICCN 3750 switch are being relocated from BRK 110 the UHB MDF. This is planned for Nov 22, during the Thanksgiving break; this should lessen the overall impact on the campus. We have connected all switches with one path to the primary distribution switch in HSB, so all the network connections will continue to function during this move. One issue will be the wireless. The secondary distribution switch has a WISM blade that controls a portion of the wireless access points (AP). We have a WISM in each of the distribution switches, and each WISM will control up to 300 AP units, we currently have 340 active APs on campus. So while the secondary distribution switch is moving there will be some AP units on campus that will not be working. The move and reactivation of the 6513 switch will take 2-3 hours.
There will be a campus wide email on Monday Nov 21, 2011 to notify the campus of the move.